find the area between y=x^2-3 and y=2x+5
Answer #1

Yani y=x²-3; 4= 2x+5 *²-3=2045 > x-ans x?-24-8-0 x-4)(x+2)50 X=-2, 4 b Areas Supperauere lower Cuare) dx (20+5)=(x²-3) de -2